OpenStudy (anonymous):

y=(3x^2-12)/(x^2-5x+6) List all the asymptotes Vertical Asymptote: Horizontal Asymptote: Oblique Asymptote: Missing Points:

OpenStudy (anonymous):

\[y=(3x^2-12)/(x^2-5x+6) \]

OpenStudy (anonymous):

y= (3x-6)/((x+3)

OpenStudy (anonymous):

ali previous question what about oblique asymptote and missing points

jhonyy9 (jhonyy9):

3x2 -12 =3(x2 -4)=3(x-2)(x+2) x2 -5x +6 =(x-3)(x-2)

OpenStudy (anonymous):

yea already at that step

OpenStudy (anonymous):

now what

OpenStudy (anonymous):

vertical asym = -3

OpenStudy (anonymous):

There is also a hole at 2

OpenStudy (anonymous):

Or missing point